Stress, elongation strain, and Young's modulus of a wire

A 100 N weight is hung from the end of a wire of cross-sectional area .010 cm^2. The wire stretches from its original length of 100.00 cm to 100.50 cm.

1. what is the stress on the wire?
2. what is the elongation strain on the wire?
3. what is the Young's modulus of the wire?

Thanks so much.
Answer #1

E=tensile stress/tensile strain

tensile stress = Force/Cross-sectional area

tensile strain = change in length / original length

tensile stress = 10^2N/1x10^-6m

tensile stress = 10^8

tensile strain = 1.0m-1.005m/1.0m

tensile strain = .005/1.0= 5x10^-3

ans= 10^8/5x10^-3

ans =0.2x10^11

answered by: olol
Answer #2

´ 102-N object is hung from the end of a wire of cross-sectional area 0.01 cm2.  The wire stretches from its original length of 200 cm to 200.5 cm